OpenStudy (anonymous):

factorize 6a2-51a+25=0; how to solve?

OpenStudy (anonymous):

factorize \[6a ^{2}-51a+25=0\] how to solve?

OpenStudy (kira_yamato):

D= 51^2 - 5(6)(25) = 2001, which is not a perfect square. Hence can't be factorized.

Directrix (directrix):

@k.m.imran Use the Quadratic Formula for this one. In case you need a review, click on the Attachment.